# CODI: Compressing Chain-of-Thought into Continuous Space via Self-Distillation

### v2

@@ Abstract
Chain-of-Thought (CoT) reasoning enhances Large Language Models (LLMs) by encouraging step-by-step reasoning in natural language. However, leveraging a latent continuous space for reasoning may offer benefits in terms of both efficiency and robustness. Prior implicit CoT methods attempt to bypass language completely by reasoning in continuous space but have consistently underperformed compared to the standard explicit CoT approach. We introduce CODI (Continuous Chain-of-Thought via Self-Distillation), a novel training framework that effectively compresses natural language CoT into continuous space. CODI jointly trains a teacher task (Explicit CoT) and a student task (Implicit CoT), distilling the reasoning ability from language into continuous space by aligning the hidden states of a designated token. Our experiments show that CODI is the first implicit CoT approach to match the performance of explicit CoT on GSM8k at the GPT-2 scale, achieving a 3.1x compression rate and outperforming the previous state-of-the-art by 28.2% in accuracy. CODI also demonstrates robustness, generalizable to complex datasets, and interpretability. These results validate that LLMs can reason effectively not only in natural language, but also in a latent continuous space.
@@ 1 Introduction
An implicit CoT algorithm replaces natural language tokens with continuous representations for reasoning as shown in Figure 1 (left). To effectively learn these representations, Pfau et al. (2024); Goyal et al. (2024) pretrain the model with additional thinking tokens from scratch. More recently, the state-of-the-art method, Coconut Hao et al. (2024) adopts a curriculum learning strategy Deng et al. (2024) that gradually replaces the initial CoT tokens with continuous thoughts. This strategy encourages continuous thoughts to behave like the removed CoT tokens. Although Coconut has greatly improved upon earlier implicit CoT methods in terms of performance  (Goyal et al., 2024; Deng et al., 2024), it lags behind CoT-SFT by a large margin as shown in Figure 1 (right). We hypothesize that this performance gap is due to forgetting across stages in the curriculum learning process Vijjini et al. (2021). This prompts us to ask: Can implicit CoT methods achieve the reasoning capability comparable to CoT-SFT while maintaining their efficiency advantages?
@@ 2 Related Work
Implicit CoT methods aim to enhance reasoning without verbalizing intermediate steps as in CoT, thereby accelerating inference speed. Theoretical work Strobl et al. (2024); Merrill and Sabharwal (2024) establishes that additional computational tokens enhance transformers’ reasoning capacity. Empirical studies Pfau et al. (2024); Goyal et al. (2024) validate these insights by training LLMs with extra dummy tokens before answering though in a limited scale and effect. Recent efforts Deng et al. (2023, 2024) distills CoT reasoning by fine-tuning. They improve over the No-CoT baseline, but fall behind CoT finetuning possibly due to discarding all intermediate tokens. Addressing this, Coconut Hao et al. (2024) reintroduces intermediate reasoning tokens via autoregressive hidden state propagation, combining curriculum learning from Deng et al. (2024). While this achieves some improvement over Deng et al. (2024), Coconut still lags behind explicit CoT, which we attribute to forgetting in curriculum learning. CODI replaces curriculum learning with a novel self-distillation framework, enabling a single-step learning process that avoids forgetting issues. Our work is also inspired by in-context compression Ge et al. (2024); Li et al. (2024c), though our work is compressing the generation instead of the existing contexts. Concurrent works Xu et al. (2025); Liu et al. (2024); Su et al. (2025) explore latent reasoning, but still rely on explicit CoT generation. Looped transformers Geiping et al. (2025a); Saunshi et al. (2025); Yu et al. (2025) also support latent reasoning, though they primarily vary in model depth without introducing. In contrast, CODI emphasizes increasing reasoning capability through additional tokens.
@@ 3.1 Overview
$\mathcal{L}=\alpha\mathcal{L}_{\text{teacher}}+\beta\mathcal{L}_{\text{student}}+\gamma\mathcal{L}_{\text{KD}},$ (1)
@@ 3.4 Self-Distillation
Rather than aligning with all tokens in the query sentence, we select a distillation token for alignment. Inspired by the recent observations (Orgad et al., 2025) that the hidden activations of the token intermediately preceding the answer, i.e., the colon (“:”) in the answer prompt “The answer is:” (as shown in Figure 2), encodes essential reasoning information. We select this token’s hidden activations, h, for distillation.
@@ 4.1 Experimental Setup
We utilize three datasets to train our models–GSM8k-Aug, GSM8k-Aug-NL, and CommonsenseQA-CoT. (1) We use the GSM8k-Aug dataset from Deng et al. (2023), which has proven effective for training implicit CoT methods Deng et al. (2024); Hao et al. (2024). This dataset extends the original GSM8k training set Cobbe et al. (2021) to 385k samples by prompting GPT-4. To facilitate implicit CoT training, all natural language interleaving within the CoT is removed, leaving only structured mathematical expressions such as “<<10÷5=2>><<2×2=4>><<6×4=24>><<10$\div$ 5=2>><<2$\times$ 2=4>><<6$\times$ 4=24>>< < 10 ÷ 5 = 2 > > < < 2 × 2 = 4 > > < < 6 × 4 = 24 > >”. (2) We also use GSM8k-Aug-NL, a version that preserves natural language explanations, to assess both the generalizability and effectiveness of our approach to compress more verbose CoTs. (3) CommonsenseQA-CoT is derived from CommonsenseQA Talmor et al. (2019), a multiple-choice QA dataset built from ConceptNet-based questions Speer et al. (2017). As it lacks CoT annotations, we generate 8.1k CoT examples using GPT-4o-mini, filtered by correctness. The 1.2k-example validation set is used for evaluation. Examples and statistics are in Appendix C.
@@ 4.4 Ablation Studies
Methods (GPT-2) Accuracy
No-CoT-SFT 19.1%
CODI 43.7%
- ind. static teacher 27.1%
         w/ multitask student 42.2%
- w/o L1 loss 24.5%
- w/ CoT last step 31.7%
- w/o Projection 42.5%
@@
To evaluate the need of self-distillation, we tested settings where the student does not share the model with the teacher (Table 2). Without learning explicit CoT generation (ind. static teacher), the model performs badly and fails to generate meaningful continuous CoTs after decoding. Adding an explicit CoT generation objective (w/ multitask student) significantly restores performance, indicating the importance of reference learning.
@@
Table 2 shows that removing the L1 loss (Equation 5) linking the teacher and student tasks (w/o L1 Loss) leads to a significant performance drop, indicating the importance of supervision from distillation. While the model performs well in CoT generation due to multitask learning, it fails to integrate this skill into continuous CoT reasoning, treating them as independent tasks rather than a unified reasoning process.
@@
Keeping the final step of the CoT chain appears to negatively impact performance, supporting our claim that it provides shortcuts. Furthermore, the projection layer of continuous thought tokens slightly enhances CODI’s effectiveness.
5 Interpretability Analysis
Figure 6: A case study illustrating CODI’s interpretability by analyzing its attended tokens and decoded tokens of each of the six latent thought tokens, $z_{1}\cdots z_{6}$. Attended tokens: these represent the top-10 tokens that the continuous thought attends to when generating the next thought/token. Some attended tokens appear in the form of ‘$z_{i}=x$’, indicating attention to the $i$-th continuous thought. Here $x$ represents the top-1 token that the latent thought maps to in vocabulary space. The model always attends to the first token in the sentence, so we remove that for better visualization. Decoded tokens: these are the top-5 words that the continuous thoughts are projected back to in vocabulary space by multiplying them with the vocabulary embeddings.
@@ 6 Conclusion
We introduced CODI, a novel paradigm for reasoning in continuous space. Our extensive experiments demonstrate CODI’s effectiveness as the new SOTA implicit CoT approach, while achieving a high compression ratio. Furthermore, CODI shows its robustness, generalisable to complex datasets, and interpretability. Future research should explore CODI’s application to more diverse and challenging tasks. A promising direction is the integration of multimodality, leveraging continuous representations for seamless modality merging. We hope this work inspires further exploration into reasoning in representations more compact and robust than language, paving the way for more efficient and versatile reasoning paradigms.
@@ 7 Limitations
Finally, while we don’t have sufficient computation resources to scale the training of CODI on larger models, a concurrent paper Geiping et al. (2025b) has demonstrated the feasibility of scaling a latent reasoning model to 3.5B parameters and 800 billion tokens with 4096 GPUs. The resulting model appears to be learning meta-strategies and abstractions for problem solving, as opposed to memorising as in existing LLMs trained on explicit CoT data. This is particularly encouraging, since not all reasoning steps can be easily verbalised (such as visual-spatial reasoning, emotional and social reasoning, and motor reasoning). While Geiping et al. (2025b) focuses on pre-training, we proposed an efficient fine-tuning approach for equipping existing pre-trained LLMs with latent reasoning capabilities.
@@ Acknowledgments
This work was supported by the UK Engineering and Physical Sciences Research Council (EPSRC) through a Turing AI Fellowship (grant no. EP/V020579/1, EP/V020579/2). ZS is supported by a PhD studentship provided by the Chinese Scholarship Council. The authors acknowledge the use of King’s Computational Research, Engineering and Technology Environment (CREATE) at King’s College London.
@@ Appendix A Implementation Details
For all experiments, we use the AdamW optimizer Loshchilov and Hutter (2019) with a cosine scheduler (without cycles) and a linear warm-up over the first 3% of steps. The effective batch size is 128. Both $\alpha$ and $\beta$ are set to 1 (Equation 1). We apply LoRA Hu et al. (2022) finetuning with a rank of 128 and an alpha value of 32, using bfloat16 precision.
@@ Appendix D CODI’s Pattern Learning
Given that CODI’s continuous thoughts can often be decoded into intermediate results, it raises a question: is CODI effectively equivalent to a GPT-2 fine-tuned on a dataset containing CODI’s decoded patterns? We created a dataset containing only intermediate results (e.g., “CoT: 20, 7, 27. Result: 9” translated from the case study in Figure 6). Additionally, since some cases of CODI show decoded operators like ‘$\times$’ and ‘--’ interleaved with intermediate results, we also create a synthetic CoT dataset that includes both operators and results (e.g., “CoT: $\times$, 20, $\times$, 7, $+$, 27. Result: 9”). As shown in Table A3, while models trained on the two synthetic datasets outperform the No-CoT-SFT baseline, they perform much worse compared to CODI, though perform on par with Coconut. These result suggest that CODI learns richer information from the teacher task through distillation than pure imitation on language-level intermediate results alone, highlighting the advantages of our training framework.
@@ Appendix E Interpretability Case Studies
Appendix F CODI Code
@@
Figure A1: CODI’s interpretability on problems involving two steps.
Figure A2: CODI’s interpretability on problems involving one step.
@@
self.prj = nn.Sequential(
nn.Linear(hidden_dim, hidden_dim),
nn.GELU(),
nn.Linear(hidden_dim, hidden_dim),
nn.LayerNorm(hidden_dim),
@@
return self.alpha*teacher_ce_loss + self.beta*student_ce_loss + self.gamma*knowledge_distillation_loss

### v3

@@ Abstract
Chain-of-Thought (CoT) reasoning enhances Large Language Models (LLMs) by encouraging step-by-step reasoning in natural language. However, leveraging a latent continuous space for reasoning may offer benefits in terms of both efficiency and robustness. Prior implicit CoT methods attempt to bypass language completely by reasoning in continuous space but have consistently underperformed compared to the standard explicit CoT approach. We introduce CODI (Continuous Chain-of-Thought via Self-Distillation), a novel training framework that effectively compresses natural language CoT into continuous space. CODI jointly trains a teacher task (Explicit CoT) and a student task (Implicit CoT), distilling the reasoning ability from language into continuous space by aligning the hidden states of a designated token. Our experiments show that CODI is the first implicit CoT approach to match the performance of explicit CoT on GSM8k at the GPT-2 scale, achieving a 3.1x compression rate and outperforming the previous state-of-the-art by 28.2% in accuracy. CODI also demonstrates robustness, generalizable to complex datasets, and interpretability. These results validate that LLMs can reason effectively not only in natural language, but also in a latent continuous space. Code is available at https://github.com/zhenyi4/codi.
@@ 1 Introduction
An implicit CoT algorithm replaces natural language tokens with continuous representations for reasoning as shown in Figure 1 (left). To effectively learn these representations, Pfau et al. (2024); Goyal et al. (2024) pretrain the model with additional thinking tokens from scratch. More recently, the state-of-the-art method, Coconut Hao et al. (2024) adopts a curriculum learning strategy Deng et al. (2024) that gradually replaces the initial CoT tokens with continuous thoughts. This strategy encourages continuous thoughts to behave like the removed CoT tokens. Although Coconut has greatly improved upon earlier implicit CoT methods in terms of performance  (Goyal et al., 2024; Deng et al., 2024), it lags behind CoT-SFT by a large margin as shown in Figure 1 (right). We hypothesize that this performance gap is due to forgetting across stages in the curriculum learning process Rao Vijjini et al. (2021). This prompts us to ask: Can implicit CoT methods achieve the reasoning capability comparable to CoT-SFT while maintaining their efficiency advantages?
@@ 2 Related Work
Implicit CoT methods aim to enhance reasoning without verbalizing intermediate steps as in CoT, thereby accelerating inference speed. Theoretical work Strobl et al. (2024); Merrill and Sabharwal (2024) establishes that additional computational tokens enhance transformers’ reasoning capacity. Empirical studies Pfau et al. (2024); Goyal et al. (2024) validate these insights by training LLMs with extra dummy tokens before answering though in a limited scale and effect. Recent efforts Deng et al. (2023, 2024) distills CoT reasoning by fine-tuning. They improve over the No-CoT baseline, but fall behind CoT finetuning possibly due to discarding all intermediate tokens. Addressing this, Coconut Hao et al. (2024) reintroduces intermediate reasoning tokens via autoregressive hidden state propagation, combining curriculum learning from Deng et al. (2024). While this achieves some improvement over Deng et al. (2024), Coconut still lags behind explicit CoT, which we attribute to forgetting in curriculum learning. CODI replaces curriculum learning with a novel self-distillation framework, enabling a single-step learning process that avoids forgetting issues. Our work is also inspired by in-context compression Ge et al. (2024); Li et al. (2025), though our work is compressing the generation instead of the existing contexts. Concurrent works Xu et al. (2025); Liu et al. (2025); Su et al. (2025) explore latent reasoning, but still rely on explicit CoT generation. Looped transformers Geiping et al. (2025); Saunshi et al. (2025); Yu et al. (2025) also support latent reasoning, though they primarily vary in model depth without introducing. In contrast, CODI emphasizes increasing reasoning capability through additional tokens.
@@ 3.1 Overview
$\mathcal{L}=\alpha\mathcal{L}_{\text{student}}+\beta\mathcal{L}_{\text{KD}}+\gamma\mathcal{L}_{\text{teacher}},$ (1)
@@ 3.4 Self-Distillation
Rather than aligning with all tokens in the query sentence, we select a distillation token for alignment. Inspired by the recent observations (Orgad et al., 2025) that the hidden activations of the token intermediately preceding the answer, i.e., the colon (“:”) in the answer prompt “The answer is:” (as shown in Figure 2), encodes essential reasoning information. We select this token’s hidden activations, h, for distillation. Alternative answer prompts and distillation tokens are also effective, and the corresponding ablation studies are reported in Appendix G.
@@ 4.1 Experimental Setup
We utilize three datasets to train our models–GSM8k-Aug, GSM8k-Aug-NL, and CommonsenseQA-CoT. (1) We use the GSM8k-Aug dataset from Deng et al. (2023), which has proven effective for training implicit CoT methods Deng et al. (2024); Hao et al. (2024). This dataset extends the original GSM8k training set Cobbe et al. (2021) to 385k samples by prompting GPT-4. To facilitate implicit CoT training, all natural language interleaving within the CoT is removed, leaving only structured mathematical expressions such as $10\div 5=2>><<2\times 2=4>><<6\times 4=24>>$”. (2) We also use GSM8k-Aug-NL, a version that preserves natural language explanations, to assess both the generalizability and effectiveness of our approach to compress more verbose CoTs. (3) CommonsenseQA-CoT is derived from CommonsenseQA Talmor et al. (2019), a multiple-choice QA dataset built from ConceptNet-based questions Speer et al. (2017). As it lacks CoT annotations, we generate 8.1k CoT examples using GPT-4o-mini, filtered by correctness. The 1.2k-example validation set is used for evaluation. Examples and statistics are in Appendix C.
@@ 4.4 Ablation Studies
! Methods (GPT-2) Accuracy No-CoT-SFT 19.1% CODI 43.7% - separate static teacher 27.1%      w/ multitask student 42.2% - w/o L1 loss 24.5% - w/ CoT last step 31.7% - w/o Projection 42.5%
@@
To evaluate the need of self-distillation, we tested settings where the student does not share the model with the teacher. As observed from Table 2, without learning explicit CoT generation (separate static teacher), the model performs badly and fails to generate meaningful continuous CoTs after decoding. Adding an explicit CoT generation objective (w/ multitask student) significantly restores performance, indicating the importance of reference learning.
@@
Table 2 also shows that removing the L1 loss (Equation 5) linking the teacher and student tasks (w/o L1 Loss) leads to a significant performance drop, indicating the importance of supervision from distillation. While the model performs well in CoT generation due to multitask learning, it fails to integrate this skill into continuous CoT reasoning, treating them as independent tasks rather than a unified reasoning process.
@@
Keeping the last step of the CoT chain appears to negatively impact performance, supporting our claim that it provides shortcuts. The projection layer of continuous thought tokens slightly enhances CODI’s effectiveness. Additional ablations on hyperparameters and the choice of distillation token are reported in Appendix F and G.
5 Further Analysis
We observe that CODI’s continuous thoughts exhibit a degree of interpretability. Notably, these patterns cannot not be trivially learned through standard token-by-token fine-tuning (see Appendix D).
Figure 6: A case study illustrating CODI’s interpretability by analyzing its attended tokens and decoded tokens of each of the six latent thought tokens, $z_{1}\cdots z_{6}$. Attended tokens: these represent the top-10 tokens that the continuous thought attends to when generating the next thought/token. Some attended tokens appear in the form of $z_{i}=x$’, indicating attention to the $i$-th continuous thought. Here $x$ represents the top-1 token that the latent thought maps to in vocabulary space. The model always attends to the first token in the sentence, so we remove that for better visualization. Decoded tokens: these are the top-5 words that the continuous thoughts are projected back to in vocabulary space by multiplying them with the vocabulary embeddings.
5.1 Interpretability Analysis
@@ 6 Conclusion
We introduced CODI, a novel paradigm for reasoning in continuous space. Our extensive experiments demonstrate CODI’s effectiveness as the new SOTA implicit CoT approach, while achieving a high compression ratio. Furthermore, CODI shows its robustness, generalisable to complex datasets, and interpretability. Future research should explore CODI’s application to more diverse and challenging tasks. We hope this work inspires further exploration into reasoning in representations more compact and robust than language, paving the way for more efficient and versatile reasoning paradigms.
@@ 7 Limitations
Finally, while we don’t have sufficient computation resources to scale the training of CODI on larger models, a concurrent paper Geiping et al. (2025) has demonstrated the feasibility of scaling a latent reasoning model to 3.5B parameters and 800 billion tokens with 4096 GPUs. The resulting model appears to be learning meta-strategies and abstractions for problem solving, as opposed to memorising as in existing LLMs trained on explicit CoT data. This is particularly encouraging, since not all reasoning steps can be easily verbalised (such as visual-spatial reasoning, emotional and social reasoning, and motor reasoning). While Geiping et al. (2025) focuses on pre-training, we proposed an efficient fine-tuning approach for equipping existing pre-trained LLMs with latent reasoning capabilities.
@@ Acknowledgments
This work was supported in part by the UK Engineering and Physical Sciences Research Council (EPSRC) (grant no. EP/V020579/1, EP/V020579/2, EP/Y003187/1, UKRI566, UKRI849). ZS is supported by a PhD studentship provided by the Chinese Scholarship Council. The authors acknowledge the use of King’s Computational Research, Engineering and Technology Environment (CREATE) at King’s College London. We thank Lin Gui for his suggestions during both the submission and rebuttal stages of this paper.
@@ Appendix A Implementation Details
For all experiments (CoT-SFT, No-CoT-SFT, and CODI) on both GSM8K and Commonsense, we use the AdamW optimizer Loshchilov and Hutter (2019) with a cosine scheduler (without cycles) and a linear warm-up over the first 3% of steps. The effective batch size is 128. Both $\alpha$ and $\beta$ are set to 1 (Equation 1). We apply LoRA Hu et al. (2022) finetuning with a rank of 128 and an alpha value of 32, using bfloat16 precision.
@@ Appendix D CODI’s Pattern Learning
Given that CODI’s continuous thoughts can often be decoded into intermediate results, it raises a question: is CODI effectively equivalent to a GPT-2 fine-tuned on a dataset containing CODI’s decoded patterns? We created a dataset containing only intermediate results (e.g., “CoT: 20, 7, 27. Result: 9” translated from the case study in Figure 6). Additionally, since some cases of CODI show decoded operators like $\times$’ and ‘-’ interleaved with intermediate results, we also create a synthetic CoT dataset that includes both operators and results (e.g., “CoT: $\times$, 20, $\times$, 7, ++, 27. Result: 9”). As shown in Table A3, while models trained on the two synthetic datasets outperform the No-CoT-SFT baseline, they perform much worse compared to CODI, though perform on par with Coconut. These result suggest that CODI learns richer information from the teacher task through distillation than pure imitation on language-level intermediate results alone, highlighting the advantages of our training framework.
@@ Appendix E Interpretability Case Studies
Appendix F Ablations on the Hyperparameter
The default settings for $\alpha$, $\beta$, and $\gamma$ from Equation 1 are 1, and we fix $\alpha=1$ for the ablations below.
$\beta$ determines the weight of the distillation loss. We find that $\beta=1$ works well for GPT-2. However, for LLaMA models, the magnitude of the distillation loss is about 10 times smaller than in GPT-2, prompting us to test larger values of $\beta$. From Table A4, increasing $\beta$ from 1 to 5 leads to a substantial accuracy improvement. Beyond $\beta$ = 5, performance plateaus, remaining stable as $\beta$ increases up to 30. Therefore, our choice of $\beta$ for LLaMA-1b is aligned with the relative scale of the distillation loss. Based on this ablation, we select $\beta$ = 20 as the default value for LLaMA-1b.
$\gamma$ determines the relative weight between the explicit CoT reasoning objective (teacher task) and the implicit CoT objective (student task) during training. Table A5 shows that a higher $\gamma$ accelerates convergence but leads to lower final performance. This likely occurs because a larger $\gamma$ encourages the model to learn more from natural language CoT reasoning (the teacher task), which serves as the main source for developing its reasoning ability and thus improves early training performance. However, since the model is ultimately evaluated on implicit CoT (the student task), which receives less emphasis during training when $\gamma$ is large, its performance on the target objective declines.
$\beta$ 1 5 10 20 30
Accuracy 46.5% 50.2% 49.1% 51.9% 51.4%
Table A4: Ablation study on $\beta$ on LLaMA-1b and GSM8k-Aug.
$\gamma$ 20 epochs 40 epochs
0.5 36.3% 38.2%
1 38.4% 43.7%
2 41.6% 41.9%
3 40.8% -
Table A5: Ablation study on $\gamma$ on GPT-2 and GSM8k-Aug. Results report accuracy (%) after training for different numbers of epochs.
Appendix G Ablations on the Choice of the Distillation Token.
ID Prompt Design Distillation Token Accuracy Within ±2×std of baseline?
1 The answer is: (baseline) : 39.0% -
2 Answer: : 38.4% Yes
3 Therefore, based on all previous calculations,
we conclude that the final answer is: : 40.2% Yes
4 The answer is is 38.1% Yes
5 We give the answer as as 40.1% Yes
6 We find the answer to be be 39.0% Yes
7 The answer is boxed{ { 38.4% Yes
Table A6: Robustness test on the answer prompt of CODI trained on GSM8k-Aug with 20 epochs.
We have conducted ablation studies to evaluate CODI’s robustness to various distillation tokens and answer prompts. As shown in Table A6, we tested a diverse set of prompts: prompts 2–3 vary the language, while prompts 4–7 focus on different distillation tokens (the last token of the prompt). To determine whether the accuracy differences are statistically significant, we follow an informal t-test approach, considering results to be significant if they fall outside the interval of ±2×std (1.8) from the baseline mean (39%), which are obtained by 5 independent runs. Our findings indicate that none of the alternative prompt designs show a statistically significant difference from the baseline, suggesting that CODI is robust to variations in both distillation tokens and answer prompt styles.
Appendix H CODI Code
@@
self.prj = $n$.Sequential(
$n$.Linear(hidden_dim, hidden_dim),
$n$.GELU(),
$n$.Linear(hidden_dim, hidden_dim),
$n$.LayerNorm(hidden_dim),
@@
return self.alpha*student_ce_loss teacher_ce_loss + self.beta*knowledge_distillation_loss + self.gamma*teacher_ce_loss
@@
Figure A2: CODI’s interpretability on problems involving two steps.
Figure A3: CODI’s interpretability on problems involving one step.